# VISTA: Knowledge-Driven Interpretable Vessel Trajectory Imputation via Large Language Models

### v1

@@ Abstract.
The Automatic Identification System provides critical information for maritime navigation and safety, yet its trajectories are often incomplete due to signal loss or deliberate tampering. Existing imputation methods emphasize trajectory recovery, paying limited attention to interpretability and failing to provide underlying knowledge that benefits downstream tasks such as anomaly detection and route planning. We propose knowledge-driven interpretable vessel trajectory imputation (VISTA), the first trajectory imputation framework that offers interpretability while simultaneously providing underlying knowledge to support downstream analysis. Specifically, we first define underlying knowledge as a combination of Structured Data-derived Knowledge (SDK) distilled from AIS data and Implicit LLM Knowledge acquired from large-scale Internet corpora. Second, to manage and leverage the SDK effectively at scale, we develop a data–knowledge–data loop that employs a Structured Data-derived Knowledge Graph for SDK extraction and knowledge-driven trajectory imputation. Third, to efficiently process large-scale AIS data, we introduce a workflow management layer that coordinates the end-to-end pipeline, enabling parallel knowledge extraction and trajectory imputation with anomaly handling and redundancy elimination. Experiments on two large AIS datasets show that VISTA is capable of state-of-the-art imputation accuracy and computational efficiency, improving over state-of-the-art baselines by 5%–94% and reducing time cost by 51%–93%, while producing interpretable knowledge cues that benefit downstream tasks. The source code and implementation details of VISTA are publicly available.
@@ 1. Introduction
The Automatic Identification System (AIS) shares the positions, identities, and other key information of vessels with nearby vessels, coastal stations and satellite, supporting tasks such as route planning, prediction, and anomaly detection (Bank, 2023). However, AIS vessel trajectories are frequently incomplete due to signal loss, equipment failure, or deliberate tampering (Łubkowski and Laskowski, 2017). To address this, trajectory imputation methods have been proposed that employ rule-based interpolation (Widyantara et al., 2023; Zaman et al., 2023; Wang et al., 2023) and advanced deep learning methods (Syed and Ahmed, 2023; Jiang et al., 2024; Zhang et al., 2021). More recently, researchers have recognized the importance of interpretable imputation, aiming to not only recover missing data but also to make the imputation process transparent and understandable to human (Alam et al., 2025; Cen et al., 2024; Fertig et al., 2025).
However, transparency by itself is often insufficient, as downstream tasks may benefit further from access to the underlying knowledge associated with each reconstructed segment: i) an estimation of the behavior pattern supported by concrete evidence, ii) a justification of the imputation result that aligns with the inferred patterns, and iii) a human-friendly explanation that conveys the deeper reasoning underlying the reconstruction of a segment. This idea is illustrated in Figure 1.
Figure 1. Example of knowledge-support imputed trajectory.
@@
AIS loses updates from a cargo ship with “under way using engine” status for about ten minutes near a busy port. The missing sub-trajectory is reconstructed as a smooth curve that connects the observed path before and after the gap. The supporting underlying knowledge can be summarized as follows:
i) Estimation of Behavior Patterns. The behavior corresponds to a decelerate-then-align pattern, frequently observed (about 68%) in AIS data for cargo vessels with queue-following intent near port.
@@
ii) Justification for the Imputation Result. The chosen imputation function (constant turn-rate with smooth speed decay) is most consistent with this pattern, accounting for over 70% of successful reconstructions for the estimated behavior pattern.
@@
iii) Underlying Cause. The behavior reflects port-entry procedures: vessels reduce speed and adjust heading to merge into inbound lanes and approach the pilot boarding line under traffic separation rules.
As illustrated to the lower-left in Figure 1, the underlying cause delivers decision-critical cues (the behavior is explained by port-entry procedures) needed for anomaly detection, sharpening the distinction between expected queue-following behavior and anomalous behavior.
This example highlights a crucial yet underexplored dimension of trajectory imputation: knowledge transfer for downstream tasks. We argue that, in real-world applications, downstream tasks generally will benefit not just from completed trajectories but also the underlying knowledge employed in the imputation process. Such knowledge allows users to enhance downstream tasks by internalizing the behavior patterns, regulatory rules and operational protocols employed during imputation, instead of depending solely on the reconstructed data.
Recent advances in large language models (LLMs) offer promising tools for addressing this challenge. LLMs possess strong capabilities in integrating contextual information, reasoning over domain knowledge, and generating human-understandable explanations (Liang et al., 2025b; Zhang et al., 2024b; Liang et al., 2025a). These properties make them well-suited for knowledge-driven trajectory imputation, where both behavioral plausibility and explanatory value are critical. Yet, applying LLMs in this setting is far from straightforward, as it raises fundamental question about how knowledge should be defined and efficiently extracted, managed, and applied. This leads to three challenges:
C1: How to define underlying knowledge that bridges interpretability and utility? A clear definition of what constitutes underlying knowledge is a prerequisite for its effective extraction, management, and utilization. The knowledge must be general enough to remain human-understandable, and to supply essential cues that can benefit downstream tasks, while also being precise enough to support trajectory imputation. To the best of our knowledge, existing methods (Musleh and Mokbel, 2023; Si et al., 2023; Yu et al., 2025) lack such a systematic formulation of knowledge, often relying on ad hoc heuristics or statistical patterns without explicitly characterizing the knowledge they employ.
C2: How to extract and utilize maritime knowledge effectively at scale? Since LLMs lack maritime domain knowledge (Yang et al., 2025), it is essential to be able to derive such knowledge directly from AIS data. However, AIS records are noisy and heterogeneous, blending meaningful behavioral patterns with irrelevant fluctuations. The key challenge is to extract reliable and interpretable knowledge from raw trajectories and to leverage it effectively to guide imputation and downstream decision-making. This requires mechanisms capable of efficiently organizing, retrieving, and applying maritime knowledge in a structured and scalable manner.
C3: How to efficiently process large-scale AIS data in a well orchestrated manner? While structured management of maritime knowledge is essential, this alone cannot support large-scale deployment. Efficient large-scale AIS processing requires workflows that integrate parallel knowledge extraction and trajectory imputation. Further, LLM inference introduces additional challenges such as stochastic responses, unstable execution times, and redundant or overlapping outputs (Ahn and Yin, 2025; Tan et al., 2025; Agrawal et al., 2024). To address these issues, workflow management must include mechanisms for fault tolerance and de-redundancy to ensure stable and efficient execution across the entire pipeline.
@@
To address these challenges, we propose VISTA, a framework for knowledge-driven interpretable vessel trajectory imputation that provides imputed trajectories together with supporting knowledge.
To address C1, we decompose underlying knowledge into two components: Structured Data-derived Knowledge (SDK) and Implicit LLM Knowledge. SDK, distilled from AIS data, captures vessel attributes, behavior patterns, imputation methods, and their statistical relations, facilitating precision and usability. Implicit LLM Knowledge, learned from large-scale Internet corpora, encodes highly compressed priors. To achieve knowledge that is generalizable for human understanding and yet precise for imputation, SDK serves as a trigger to retrieve and align relevant priors from the LLM, and both are integrated into a unified natural-language format.
To address C2, we establish a data–knowledge–data loop with a Structured Data-derived Knowledge Graph (SD-KG) at its core. The SD-KG serves as a structured and compact representation that organizes vessel attributes, behavior patterns, and imputation methods extracted from AIS data. When going from data to knowledge, VISTA transforms raw AIS data into structured knowledge units by eliminating noise, abstracting behavior patterns, and generating imputation methods, and it then integrates them into the SD-KG. When going from knowledge to data, VISTA leverages the accumulated knowledge in the SD-KG to reconstruct missing trajectories by estimating behaviors and selecting suitable imputation methods guided by evidence. The reconstructed trajectories are then augmented with explanatory context derived from the SD-KG, the goal being to achieve accuracy, interpretability, and applicability to downstream maritime tasks.
To address C3, we design a workflow management layer that enables efficient and reliable processing of large-scale AIS data. It manages the end-to-end pipeline from knowledge extraction to trajectory imputation through coordinated task scheduling, anomaly handling, and redundancy control. This design addresses the inherent instability and inefficiency of LLM inference by introducing mechanisms for parallel execution, fault tolerance, and quality validation, ensuring that large numbers of inference jobs can be executed consistently and efficiently.
The main contributions are summarized as follows:
  • We propose VISTA, a knowledge-driven and interpretable trajectory imputation framework that integrates structured knowledge from AIS data with implicit knowledge from LLMs to enable both reconstruction and explanation.
  • We define underlying knowledge as the combination of Structured Data-derived Knowledge and Implicit LLM Knowledge, and organize it in a Structured Data-derived Knowledge Graph with algorithms for construction, maintenance, and utilization.
@@
  • We design a workflow management layer that coordinates knowledge extraction and trajectory imputation with parallel scheduling, anomaly handling, and redundancy control, enabling efficient and reliable processing of large-scale AIS data.
@@
  • Experiments on two AIS datasets show that VISTA achieves state-of-the-art accuracy, outperforming existing baselines by 5%–94% and reducing inference time by 51%–93%, while producing knowledge cues for downstream applications.
@@
The rest of the paper is structured as follows: Section 2 reviews related work. Section 3 covers notation, the definition of underlying knowledge, and the problem formulation. Section 4 presents VISTA. Section 5 reports experimental results. Section 6 concludes the paper and outlines research directions.
@@ 2. Related Work
Table 1 summarizes four kinds of existing trajectory imputation methodologies across three key dimensions: generalization, interpretability, and the ability to provide explicit knowledge. Below we discuss the four major categories in detail.
Rule-based Trajectory Imputation Methods. Prior to the advent of data-driven models, trajectory imputation relied on mathematical rules and physical principles. Classical interpolation methods such as Linear and Cubic Spline Interpolation (Widyantara et al., 2023; Zaman et al., 2023) provide simple, interpretable solutions but ignore geographic or navigational constraints. Extensions like Cubic Hermite Splines incorporate velocity and heading to better capture vessel dynamics, while hybrid methods combine linear and spline interpolation for improved accuracy (Widyantara et al., 2023). Kinematic state estimation methods (e.g., EKF/UKF with CTRA or CMM models (Wang et al., 2023)) enforce consistency with motion laws but remain sensitive to abrupt maneuvers and data sparsity. Overall, rule-based methods are computationally efficient and interpretable, but they lack contextual awareness and generalization to complex maritime environments (Li et al., 2023; Bai et al., 2024; Fu et al., 2017).
Deep Learning-based Trajectory Imputation Methods. With the availability of large-scale AIS data, deep learning has become the dominant paradigm for trajectory imputation. Early works relied on recurrent models such as RNNs, LSTMs, and GRUs, which capture temporal dependencies and overcome the vanishing gradient problem, often combined with clustering strategies to specialize models for distinct shipping routes (Shin and Yang, 2024; Yang et al., 2022; Zhang et al., 2020). Later studies integrated Convolutional Neural Networks with recurrent networks to jointly model local spatial features and temporal evolution (Syed and Ahmed, 2023), while Transformer-based architectures (Wang et al., 2024) now represent the state-of-the-art. Their self-attention mechanism captures long-range dependencies critical for understanding navigational intent. In parallel, Graph Neural Networks (GNNs) (Jiang et al., 2024) have been applied to encode multi-vessel interactions by representing vessels as nodes and their spatial relations as edges, enabling cooperative trajectory prediction in congested waters. Generative methods, particularly GANs (Zhang et al., 2021), Diffusion Models (Zhang et al., 2025; Tashiro et al., 2021; Chen et al., 2024), provide another line of research by learning realistic distributions of vessel motion, producing more diverse and plausible imputations than conventional regression-based methods.
While deep learning methods achieve superior accuracy and strong generalization by directly learning non-linear spatio-temporal patterns, they suffer from poor interpretability. Their decision-making process remains opaque, making it difficult to assess whether outputs are grounded in navigational knowledge or merely reflect statistical correlations. Moreover, these methods do not provide explicit knowledge that can be reused or audited, which limits trust in safety-critical applications (Zhang et al., 2024a).
Table 1. Comparison of existing imputation methods.
Methodology Class Generalization Interpretability Explicit Knowledge
Rule-based Trajectory Imputation No Yes No
Deep Learning-based Trajectory Imputation Yes No No
LLM-based Trajectory Imputation Yes Yes No
Interpretable Trajectory Imputation Yes Yes No
LLM-based Trajectory Imputation Methods. A recent paradigm explores reframing trajectory imputation as a language modeling problem by treating a sequence of GPS points as analogous to words in a sentence. Through spatial discretization into a finite “location vocabulary,” vessel movement can be modeled with a “physical grammar,” enabling the use of Transformer-based architectures such as BERT for trajectory infilling tasks. Representative systems include KAMEL (Musleh and Mokbel, 2023) , which introduces spatially-aware tokenization and multi-point generation to address the limitations of vanilla LLMs, and TrajBERT (Si et al., 2023) , which incorporates spatio-temporal refinement strategies for sparse AIS data. More recent frameworks (e.g., MAKER (Yu et al., 2025) , AIS-LLM (Park et al., 2025) ) extend beyond single-task imputation by fusing trajectories with textual prompts or unifying tasks such as prediction, anomaly detection, and risk assessment, often producing natural language explanations to enhance user trust.
Despite these advances, LLM-based methods still face several fundamental challenges. Tokenizing continuous spatio-temporal data remains non-trivial, generated trajectories are often not physically grounded, and performance can be sensitive to prompt design (Xu et al., 2025). While these models offer stronger process interpretability than deep learning-based methods, they still lack mechanisms to provide explicit, structured knowledge that supports transparent reasoning and consistent integration with downstream tasks.
Interpretable Trajectory Imputation Methods. To overcome the opacity of black-box models in safety-critical maritime systems, interpretable imputation methods aim to balance predictive accuracy with transparency. A prominent direction is the use of Physics-Informed Neural Networks (Alam et al., 2025) , which embed kinematic equations into the loss function so that outputs remain consistent with physical laws. This improves generalization and robustness, particularly under sparse or noisy AIS data. Another line of work (Cen et al., 2024) leverages attention mechanisms to visualize which historical trajectory points influence predictions, providing procedural transparency and aligning model reasoning with domain expertise. Hybrid methods (Fertig et al., 2025) combine neural networks with explicit kinematic models by predicting intermediate physical variables (e.g., acceleration), ensuring that the final trajectory remains physically feasible.
These methods can be seen as either constraining outputs to a “glass box” of physical laws or offering whiteboard-style insights into the model’s internal focus. While such designs enhance both interpretability and generalization compared to pure deep learning, they still lack the ability to provide explicit knowledge that downstream tasks and operators can directly reuse.

### v2

# arXiv:2601.06940v2[cs.DB] 27 Feb 2026

@@ Abstract.
Repairing incomplete trajectory data is essential for downstream spatio-temporal applications. Yet, existing repair methods focus solely on reconstruction without documenting the reasoning behind repair decisions, undermining trust in safety-critical applications where repaired trajectories affect operational decisions, such as in maritime anomaly detection and route planning. We introduce repair provenance—structured, queryable metadata that documents the full reasoning chain behind each repair—which transforms imputation from pure data recovery into a task that supports downstream decision-making. We propose VISTA (knowledge-driven interpretable vessel trajectory imputation), a framework that reliably equips repaired trajectories with repair provenance by grounding LLM reasoning in data-verified knowledge. Specifically, we formalize Structured Data-derived Knowledge (SDK), a knowledge model whose data-verifiable components can be validated against real data and used to anchor and constrain LLM-generated explanations. We organize SDK in a Structured Data-derived Knowledge Graph (SD-KG) and establish a data–knowledge–data loop for extraction, validation, and incremental maintenance over large-scale AIS data. A workflow management layer with parallel scheduling, fault tolerance, and redundancy control ensures consistent and efficient end-to-end processing. Experiments on two large-scale AIS datasets show that VISTA achieves state-of-the-art accuracy, improving over baselines by 5%–91% and reducing inference time by 51%–93%, while producing repair provenance, whose interpretability is further validated through a case study and an interactive demo system.
@@ 1. Introduction
Repairing incomplete trajectory data is essential for downstream applications such as route planning, traffic monitoring, and anomaly detection (Fan et al., 2008; Rekatsinas et al., 2017; Asian Development Bank, 2023; Song et al., 2015; Rezig et al., 2021). While existing repair methods—whether rule-based (Widyantara et al., 2023; Wang et al., 2023), learning-based (Jiang et al., 2024; Syed and Ahmed, 2023; Liu et al., 2025a), or LLM-driven (Narayan et al., 2022; Musleh and Mokbel, 2023)—differ in transparency, none produces structured, queryable metadata that documents the reasoning behind repair decisions. In safety-critical domains such as maritime safety via the Automatic Identification System (AIS) (Tu et al., 2018), where repaired data affects operational decisions, the absence of such metadata undermines the trustworthiness of the entire analytical pipeline (Cheney et al., 2009).
Classical data provenance (Buneman et al., 2001; Cheney et al., 2009) tracks how query results are derived from source tuples, but does not capture the reasoning behind repair decisions. To bridge this gap, we introduce repair provenance, structured and queryable metadata that records the full reasoning chain behind each repaired value through three components: i) what behavior pattern is estimated for the missing segment, supported by statistical evidence; ii) how the segment is imputed, justified by the inferred pattern; and iii) why the behavior occurred, grounded in a human-readable causal explanation.
Figure 1. An imputed trajectory with repair provenance. Upper-left: observed (purple) and imputed (red) segments near a port. Right: provenance components—behavior pattern, imputation function, and causal explanation. Lower-left: provenance-supported anomaly detection.
@@
A cargo ship’s AIS messages (status: “under way using engine”) are lost for about ten minutes near a busy port. The missing sub-trajectory (red curve in Figure 1) is reconstructed as a smooth curve between the observed segments (purple points). The accompanying repair provenance (right-side panels) encompasses three components:
i) What behavior pattern is estimated. The estimated pattern is a deceleration from 7.2 to 5.0kn with a smooth ${\sim}80^{\circ}$ port-side alignment, heading fluctuation $2^{\circ}$, and duration ${\approx}600$s, indicating queue-following intent. In historical AIS data, cargo vessels approaching port entrances exhibit this pattern in 68% of comparable gaps.
@@
ii) How the segment is imputed. The selected imputation function, constant turn-rate with smooth speed decay (CTR+SD), is statistically the most consistent with the estimated pattern, accounting for over 70% of successful reconstructions under this pattern.
@@
iii) Why the behavior occurred. The behavior reflects port-entry procedures: vessels reduce speed and adjust heading to merge into inbound lanes and approach the pilot boarding line under traffic separation rules. This causal explanation distinguishes routine port-entry maneuvers from anomalous behaviors such as evasive steering or unauthorized anchoring.
As illustrated in the lower-left panel, this provenance directly supports downstream anomaly detection, enabling the operator to judge the vessel as non-anomalous.
This illustrates how repair provenance can transform trajectory imputation from a pure data reconstruction task to one that better supports downstream decision-making. Large language models (LLMs) (Liang et al., 2025b; Zhang et al., 2024; Liang et al., 2025a), with their capacity for contextual reasoning, knowledge integration, and text generation, offer the first technical means to produce such provenance automatically, yet leveraging them is challenging.
C1: How to define a knowledge model that grounds and constrains LLM reasoning? LLM outputs are unreliable—they may hallucinate behavior patterns, generate invalid imputation functions, or fabricate statistical evidence (Shankar and others, 2024; Agrawal et al., 2024b). To produce trustworthy repair provenance, a structured knowledge model is needed whose data-verifiable components—behavior patterns, imputation functions, and their statistical associations—can be validated against real data and used to anchor and constrain LLM-generated explanations. No existing formalism provides this: classical provenance models (Buneman et al., 2001; Cheney et al., 2009; Green et al., 2007) track query derivations but not repair reasoning, while data repair systems (Rekatsinas et al., 2017; Geerts et al., 2013; Rammelaere and Geerts, 2018) either lack explanations or produce unverified, unstructured annotations.
C2: How to construct and maintain a data-verified knowledge base? The knowledge model above provides a structural blueprint, but its components must be populated with empirically grounded content—behavior patterns statistically supported by observed data, imputation functions that execute correctly on real trajectories, and associations backed by verified evidence (Yang et al., 2025; Fernandez and others, 2023). This requires extracting structured knowledge from massive, noisy trajectory records, validating each component through execution and statistical testing, and updating the knowledge base as new data arrives (Hogan et al., 2021; Liang et al., 2025a). Only then can it reliably ground and constrain LLM reasoning for provenance generation.
C3: How to impute trajectories with repair provenance reliably at scale? The full pipeline—from knowledge extraction through trajectory imputation to provenance generation—involves numerous LLM inference calls that may vary across runs, may time out, or may produce redundant artifacts (Ahn and Yin, 2025; Tan et al., 2025; Agrawal et al., 2024a). Moreover, the steps are interdependent: a failed pattern extraction affects subsequent imputation and provenance steps. An execution framework must therefore ensure coordinated parallel scheduling, fault tolerance, and redundancy control to achieve consistent end-to-end processing.
@@
To address these challenges, we propose VISTA, a knowledge-driven interpretable vessel trajectory imputation framework that equips repaired trajectories with structured repair provenance.
To address C1, we propose Structured Data-derived Knowledge (SDK), a knowledge model that organizes data-verifiable knowledge, vessel attributes, behavior patterns, imputation functions, and their statistical associations, in a relational schema suitable for storage, querying, and validation. Validated SDK serves as grounded context for steering LLM reasoning towards correct causal explanations rather than hallucinated ones. Together, SDK-derived facts and LLM-generated explanations form complete repair provenance records.
To address C2, we establish a data–knowledge–data loop with a Structured Data-derived Knowledge Graph (SD-KG) at its core. Going from data to knowledge, VISTA extracts knowledge units from raw trajectory records and validates each unit through execution and statistical testing before admitting it into the SD-KG. Going from knowledge to data, validated SD-KG guides trajectory imputation and provenance generation. The repaired data is then fed back into the SD-KG, closing the loop and incrementally enriching the knowledge base with every repaired trajectory.
To address C3, we design a workflow management layer that coordinates the full end-to-end pipeline. It employs parallel task scheduling, fault tolerance, and redundancy control to handle the instability inherent in large-scale LLM inference. This ensures that failures at individual steps do not cascade and that the end-to-end process remains consistent and efficient.
To validate the generated repair provenance, we provide a case study that qualitatively examines its interpretability and develop CLEAR (Liu et al., 2025b), an interactive demo system that enables non-expert users to explore, complete, and annotate vessel trajectories through a knowledge-graph-driven interface. The main contributions are summarized as follows:
  • We propose VISTA, a knowledge-driven interpretable vessel trajectory imputation framework that reliably equips repaired trajectories with structured repair provenance at scale by grounding LLM reasoning in data-verified knowledge.
  • We formalize Structured Data-derived Knowledge and organize it in a Structured Data-derived Knowledge Graph, establishing a data–knowledge–data loop for extraction, validation, incremental maintenance, and context-aware querying.
@@
  • We design a workflow management layer featuring parallel scheduling, fault tolerance, and redundancy control to enable consistent and efficient processing of large-scale AIS data.
@@
  • Experiments on two AIS datasets show that VISTA is capable of state-of-the-art accuracy, outperforming existing baselines by 5%–91% and reducing inference time by 51%–93%, while producing repair provenance.
@@
The rest of the paper is structured as follows: Section 2 reviews related work. Section 3 covers notation, the knowledge model, and the problem formulation. Section 4 presents VISTA. Section 5 reports experimental results, and Section 6 concludes the paper.
@@ 2. Related Work
Leveraging LLMs to produce repair provenance for trajectory imputation touches three areas: trajectory imputation, data provenance and repair explanation, and knowledge-enhanced LLM reasoning.
Trajectory Imputation. Trajectory imputation has been studied using four approaches, each advancing reconstruction accuracy while leaving repair decisions undocumented. Rule-based methods repair data by enforcing declarative constraints—functional dependencies (FDs), such as conditional FDs (CFDs) (Fan et al., 2008), and denial constraints (Chu et al., 2013)—or domain-specific speed and acceleration bounds (Song et al., 2015), with scalable systems such as Horizon (Rezig et al., 2021) enabling FD-based cleaning over large datasets. These methods require manual rule specification and do not record why a particular repair value was chosen over alternatives. Interpolation-based methods fit predefined mathematical models to observed data, such as linear and spline interpolation (Widyantara et al., 2023; Zaman et al., 2023). These methods are used widely as preprocessing steps in trajectory analysis pipelines due to their simplicity and low computational cost. However, they assume smooth dynamics between the observed points and offer limited adaptability to complex movement patterns such as abrupt turns or speed changes. Deep learning-based methods learn spatio-temporal patterns directly from large-scale data. General-purpose methods such as GAIN (Yoon et al., 2018), SAITS (Du et al., 2023), CSDI (Tashiro et al., 2021), and DiffPuter (Zhang et al., 2025) handle missing multivariate data through adversarial training, self-attention, or diffusion models, while BERT-based methods such as KAMEL (Musleh and Mokbel, 2023) and TrajBERT (Si et al., 2024) adapt masked-token pre-training to spatial sequences. These methods achieve state-of-the-art reconstruction accuracy but remain black boxes with no transparency into individual repair decisions. LLM-based methods leverage the contextual reasoning and text generation capabilities of LLMs for data management tasks. For data cleaning, foundation models have been evaluated on imputation and error detection (Narayan et al., 2022), and RetClean (Naeem et al., 2024) combines retrieval from data lakes with LLM inference to repair dirty data. In the maritime domain, LLMs have been applied to related trajectory tasks such as prediction and anomaly detection (Yu et al., 2025; Park et al., 2025), but not yet to imputation with provenance. While LLM-based methods can produce free-text justifications alongside their outputs, such explanations are post-hoc rationalizations rather than faithful accounts of the repair process (Xu et al., 2025; Li et al., 2024), and they are neither structured nor queryable. Across all four approaches, no method produces structured repair provenance alongside the imputed values.
Data Provenance and Repair Explanation. Classical data provenance tracks how query results derive from source tuples. Thus, why- and where-provenance (Cheney et al., 2009; Buneman et al., 2001) identify contributing source tuples and pinpoint copied attribute values, respectively, while how-provenance (Green et al., 2007) captures computation structure via semiring annotations. Data lineage (Cui et al., 2000) and provenance-aware systems such as Perm (Glavic and Alonso, 2009) further support debugging and auditing through query rewriting. Such provenance is structured and queryable via standard SQL. However, classical provenance tracks data derivation—how query results are computed from existing data—rather than data repair reasoning—why a missing value was replaced with a specific alternative over others.
Several systems aim to justify individual repair decisions. NADEEF (Dallachiesa et al., 2013) stores violation and repair metadata in relational tables, recording which rule triggered each change, but does not capture why a particular replacement value was chosen over alternatives. LLUNATIC (Geerts et al., 2013) derives repairs through chase-based reasoning with formal guarantees. Yet, its justification is implicit in the chase trace rather than materialized as queryable metadata. HoloClean (Rekatsinas et al., 2017) outputs per-value confidence scores through probabilistic inference, but these scores reflect global model weights rather than per-value signal attributions. XPlode (Rammelaere and Geerts, 2018) identifies the minimal CFD set, justifying a group of repairs, producing constraint-level, rather than per-value, explanations. Complementary efforts address related goals: CAPE (Miao et al., 2019) explains query outliers via pattern-based counterbalances; counterfactual methods (Galhotra et al., 2021) and Gopher (Pradhan et al., 2022) reveal ML decision boundaries and bias sources; self-supervised methods (Peng et al., 2022) learn interpretable repair rules; and in-database imputation (Perini and Nikolic, 2024) integrates MICE within a DBMS.
In summary, classical provenance is structured and queryable but does not address repair reasoning. Existing repair explanation methods provide partial justification—audit logs, confidence scores, or constraint references—but none produces a unified, queryable record of the full reasoning chain behind each repair.
Knowledge-Enhanced LLM Reasoning. Knowledge graphs provide structured, queryable representations of domain facts (Hogan et al., 2021) that have recently been leveraged to ground LLM reasoning and reduce hallucination (Pan et al., 2024; Agrawal et al., 2024b; Yang et al., 2025). Pan et al. (Pan et al., 2024) provide a roadmap categorizing approaches into KG-enhanced LLMs, LLM-augmented KGs, and synergized LLM+KG systems. KAG (Liang et al., 2025a) combines logical-form-guided retrieval with vector search to enhance LLM reasoning in professional domains. On the retrieval side, retrieval-augmented generation (RAG) (Lewis et al., 2020) anchors LLM outputs in external evidence, and graph-structured extensions (Peng et al., 2025) further incorporate relational knowledge into the retrieval process.
However, existing methods are designed primarily for open-domain question answering or text generation; the knowledge they retrieve is typically accepted without validation against the data. At scale, this absence of validation is compounded by LLM-specific reliability challenges, prompt inconsistency (Ahn and Yin, 2025), self-consistent errors (Tan et al., 2025), and resource contention (Agrawal et al., 2024a), that can degrade output quality. No existing studies validate knowledge against the data before using it to guide LLM reasoning; nor do they provide pipeline-level mechanisms to ensure reliable execution at scale.